# Thermal spectrometer for superconducting circuits

### v1

@@ Conclusion
In summary, we have demonstrated operation of an on-chip bolometric spectrometer to characterize a superconducting resonator. The measurement is done by a simple dc measurement setup. The bolometer operates by absorbing the energy decay of the resonator, leading to temperature rise in the bolometer, which is detected by measuring the dc voltage change $V_{\text{th}}$ acrross a pair of NIS thermometer junctions. The resonance frequency and the lineshape (quality factor) of the resonator can be determined based on the temperature change. We have demonstrated additional advantages of the bolometer compared to standard rf measurement in the high frequency range: the bandwidth of the bolometer exceeds that of a standard rf spectrometer. Furthermore, we found a calibration-free regime where the measured voltage $V_{\text{th}}$ follows Lorentzian spectrum. The thermal spectrometer could then be used in the single-photon regime to perform both one-tone and two-tone spectroscopy of a qubit.
@@ Methods
The fabrication of the device is done in a multistage process on a $675~{}\mathrm{\mu m}$-thick and highly resistive silicon (Si) substrate, resulting to device shown in Fig. 2a-c. The fabrication consists of there main steps: (1) fabricating niobium (Nb) structures (resonator, feedline, ground plane, and pads), (2) fabrictaing flux qubit made of three junctions of superconductor–insulator–superconductor (SIS) with aluminium (Al) film, (3) fabricating bolometer consisting absorber and thermometer. A $40~{}\mathrm{nm}$-thick Al2O3 layer is deposited onto a silicon substrate using atomic layer deposition, followed by a deposition of a $200~{}\mathrm{nm}$-thick Nb film using DC magnetron sputtering. Positive electron beam resist, AR-P6200.13, is spin-coated with a speed of 6000 rpm for 60 s, and is post-baked for 9 minutes at 160C, which is then patterned by electron beam lithography (EBL) and etched by reactive ion etching. A shadow mask defined by EBL on a 1 $\mathrm{\mu m}$-thick poly(methyl-metacrylate)/copolymer resist bilayer is used to fabricate the flux qubit made of Al film shunting the resonator at the shorted end [30]. Before the deposition of Al film, the Nb surface is cleaned in-situ by Ar ion plasma milling for 60 s. The bolometer structure is fabricated with three angle deposition technique. To have a clean contact between Nb film and Al film, the Nb surface is cleaned in-situ by Ar ion plasma milling for 45 s, followed by deposition at +40 of 20 nm-thick Al lead. The Al lead is oxidized at 2.5 mbar pressure for 2 minutes. After that, 3 nm-thick Al buffer layer is deposited at -6.5, followed by deposition of 30 nm-thick Cu film at -6.5. Finally, 90 nm-thick Al film is deposited at +20 on top of the edge of Cu film to connect the Cu to the Nb capacitor. Finally, after liftoff in hot acetone (52 degrees for  30 minutes) and cleaing with isopropyl alcohol, the substrate is cut by an automatic dicing-saw machine to the size $7\times 7$ mm and wire-bonded to an cooper rf-dc holder for the low-temperature characterization.
@@
Measurements are performed in a cryogen-free dilution refrigerator at temperature of 30 mK with the setup shown in Fig. 2d. To obtain scattering data $S_{21}$, using a VNA, a probe microwave tone is supplied to the input feedline through a 90 dB of attenuation distributed at the various temperature stages of the fridge. The output probe signal is then passed through two cryogenic circulators, before being amplified first by a 40 dB cryogenic amplifier and then by a 40 dB room-temperature amplifier. The volatge $V_{\text{th}}$ is measured by a dc voltmeter through thermocoax line. The device is mounted in a tight cooper holder and covered by an Al shield.

### v2

@@ Conclusion
In summary, we have demonstrated operation of an on-chip bolometric spectrometer to characterize a superconducting resonator. The measurement is done by a simple DC measurement setup. The bolometer operates by absorbing the energy decay of the resonator, leading to temperature rise in the bolometer, which is detected by measuring the DC voltage change $V_{\text{th}}$ acrross a pair of NIS thermometer junctions. The resonance frequency and the lineshape (quality factor) of the resonator can be determined based on the temperature change. We have demonstrated additional advantages of the bolometer compared to standard RF measurement in the high frequency range: the bandwidth of the bolometer exceeds that of a standard RF spectrometer. Furthermore, we found a calibration-free regime where the measured voltage $V_{\text{th}}$ follows Lorentzian spectrum. The thermal spectrometer could then be used in the single-photon regime to perform both one-tone and two-tone spectroscopy of a qubit (see Supplementary Information V for a preliminary result).
@@ Methods
The fabrication of the device is done in a multistage process on a $675~{}\mathrm{\mu m}$-thick and highly resistive silicon (Si) substrate, resulting to device shown in Fig. 2a-c. The fabrication consists of three main steps: (1) fabricating niobium (Nb) structures (resonator, feedline, ground plane, and pads), (2) fabrictaing flux qubit made of three junctions of superconductor–insulator–superconductor (SIS) with aluminium (Al) film, (3) fabricating bolometer consisting absorber and thermometer. A $40~{}\mathrm{nm}$-thick AlOx layer is deposited onto a silicon substrate using atomic layer deposition, followed by a deposition of a $200~{}\mathrm{nm}$-thick Nb film using DC magnetron sputtering. Positive electron beam resist, AR-P6200.13, is spin-coated with a speed of 6000 rpm for 60 s, and is post-baked for 9 minutes at 160C, which is then patterned by electron beam lithography (EBL) and etched by reactive ion etching. A shadow mask defined by EBL on a 1 $\mathrm{\mu m}$-thick poly(methyl-metacrylate)/copolymer resist bilayer is used to fabricate the flux qubit made of Al film shunting the resonator at the shorted end [30]. Before the deposition of Al film, the Nb surface is cleaned in-situ by Ar ion plasma milling for 60 s. The bolometer structure is fabricated with three angle deposition technique. To have a clean contact between Nb film and Al film, the Nb surface is cleaned in-situ by Ar ion plasma milling for 45 s, followed by deposition at +40 of 20 nm-thick Al lead. The Al lead is oxidized at 2.5 mbar pressure for 2 minutes. After that, 3 nm-thick Al buffer layer is deposited at -6.5, followed by deposition of 30 nm-thick Cu film at -6.5. Finally, 90 nm-thick Al film is deposited at +20 on top of the edge of Cu film to connect the Cu to the Nb capacitor. Finally, after liftoff in hot acetone (52 degrees for  30 minutes) and cleaning with isopropyl alcohol, the substrate is cut by an automatic dicing-saw machine to the size $7\times 7$ mm and wire-bonded to an copper RF-DC holder for the low-temperature characterization.
@@
Measurements are performed in a cryogen-free dilution refrigerator at temperature of 50 mK with the setup shown in Fig. 2d. To obtain scattering data $S_{21}$, using a VNA, a probe microwave tone is supplied to the input feedline through a 90 dB of attenuation distributed at the various temperature stages of the fridge. The output probe signal is then passed through two cryogenic circulators, before being amplified first by a 40 dB cryogenic amplifier and then by a 40 dB room-temperature amplifier. The volatge $V_{\text{th}}$ is measured by a DC voltmeter through thermocoax line. The device is mounted in a tight cooper holder and covered by an Al shield.
Data availability
The data that support the plots within this article are available from the corresponding author upon reasonable request.
Code availability
The codes that support the findings of this study are available from the corresponding author upon request.
@@ Acknowledgements
Author contributions
C.D.S, Y.C.C, and J.P.P conceived the idea of the experiment. C.D.S designed the device with inputs from Y.C.C and R.U. C.D.S fabricated the device supported by J.T.P and R.U. C.D.S performed the experiments supported by Y.C.C and J.T.P. J.P.P, B.K, I.K.M performed theoretical analysis and wrote supplementary information I, II, and IV. The data analysis was conducted by C.D.S, Y.C.C, and A.S.S with inputs from all the authors. The manuscript was written by C.D.S with contributions from all the authors. The figures were prepared by C.D.S and B.K.
Competing interests
The authors declare no competing interests.